### v1

@@ 3.4 Recovery performance analysis of confined OMP
where $\mathbb{P}\{\nu^{(k)}=\upsilon\}$ is given in Lemma 3.1.
Proof 3.9.
See Appendix E.
4 Experimental results
This section presents the experimental results that demonstrate the advantage of the confined OMP algorithm in terms of recovery performance and complexity. Furthermore, we use the lower bound on the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ to optimize the column degree $d$. The experimental results were obtained by MATLAB R2023a on a desktop computer with Intel(R) Core(TM) i7-11700 CPU @ 2.50 GHz. In the following simulations, we generate a $K$-sparse signal whose support is chosen at random. In addition, we consider two types of $K$-sparse signals: (a) Gaussian signals and (b) flat signals. The non-zero elements of the Gaussian signal are independently and randomly drawn from a standard Gaussian distribution. Furthermore, the support of the flat signal is randomly chosen and the nonzero elements are set to 1.
Figure 3: Recovery performance for $K$-sparse (a) Gaussian signals and (b) flat signals versus the sparsity $K$.
Figure 4: (a) Average number of inner-product operations in the identification step and (b) average CPU times of different greedy algorithms for Gaussian signals with different $K$.
4.1 Recovery performances for different sparsity $K$
We first show the efficiency and recovery performance of the proposed confined OMP algorithm for different sparsity $K$. In Fig. 3 and Fig. 4, each column of the measurement matrix $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ is chosen independently and uniformly among the vectors with $10$ ones.
In Fig. 3, we can see that the confined OMP algorithm achieves performance gains for both Gaussian and flat signals. In particular, the performance gain for the flat signal is appreciable. It is reasonable because when the sparsity $K$ is relative small, the size of $\Gamma$ is not as large, which enables the confined OMP algorithm to screen out more interference.
Figure 5: Recovery performance for $4$-sparse (a) Gaussian signals and (b) flat signals versus measurement $m$.
Figure 6: (a) Average number of inner-product operations in the identification step and (b) average CPU times of different greedy algorithms for Gaussian signals with different measurements $m$.
Fig. 4 shows the average inner-product operations in the identification step and average CPU times of different greedy algorithms for Gaussian signals with different $K$. From Fig. 4(a), we observe that the number of inner-product operations of OMP algorithm increases linearly with the increase of $K$. The reason is that OMP algorithm requires $Kn$ inner-product operations in the identification step to recover a $K$-sparse signal. For the confined OMP algorithm, however, the identification is eliminated if $|\Gamma|=K$. Otherwise, $K|\Gamma|$ inner-product operations are required to recover a $K$-sparse signal. As shown in Fig. 2, the expectation of $|\Gamma|$ is much smaller than $n$ for a relatively small $K$. Thus, in Fig. 4(a), the confined OMP algorithm achieves a significant reduction in number of inner-product operations for a relatively small $K$. The average CPU time shown in Fig. 4(b) relates to the number of inner-product operations. Their trends of curves shown in Fig. 4(a) and (b) are almost the same. In particular, the confined OMP algorithm achieves a reduction of about $85\%$ in average CPU time to recover a $6$-sparse Gaussian signal when compared to that of OMP algorithm. Furthermore, we also present the average CPU time of the batch OMP algorithm proposed in [32] for comparison. It can be seen that the complexity reduction of the confined OMP algorithm is much greater than that of the batch OMP algorithm for a relatively small $K$.
Remark: The batch OMP algorithm is an efficient implementation of the OMP algorithm. Their recovery performances are the same. Thus, we omit the performance comparison between the batch OMP algorithm and the proposed algorithm. More details about the batch OMP algorithm can be found in [32].
4.2 Recovery performances for different measurements $m$
We further show the efficiency and recovery performance of the confined OMP algorithm for different measurements $m$. In Fig. 5 and Fig. 6, each column of the measurement matrix $\boldsymbol{A}\in\{0,1\}^{m\times 256}$ is chosen independently and uniformly among the vectors with $10$ ones. The sparsity $K$ is set to $4$.
In Fig. 5, the recovery performances of confined OMP algorithm are better than those of OMP algorithm, but performance gains become smaller as $m$ increases. However, we observe in Fig. 6 that the number of inner-product operations for the confined OMP algorithm tends to 0 as $m$ increases, resulting in a reduction of about $86\%$ in average CPU time. These observations indicate that the complexity of the confined OMP algorithm is mainly contributed by the complexity of solving least-squares problems when $m\geq 100$.
4.3 Lower bounds
The lower bounds on the probability of $\mathbb{S}_{\mathrm{omp}}$ and $\mathbb{S}_{\mathrm{comp}}$ can be obtained by Eq. 9 and Eq. 10, respectively. In experimental tests, the column degree $d$ of each chosen $\boldsymbol{A}$ satisfies $\log m<d\leq m/2$ for a sufficiently large $m$. In this case, the probability $\mathbb{P}\left\{\det\left(\boldsymbol{A}_{\Lambda^{(k)}}^{T}\boldsymbol{A}_{\Lambda^{(k)}}\right)\neq 0\right\}$ is approximately equal to 1. Thus, the effect of the second term for both Eq. 9 and Eq. 10 is negligible.
The lower bounds on $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$ and $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ are shown in Fig. 3 and Fig. 5. It can be seen that the lower bound on $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ is much tighter than that on $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$. However, for both OMP algorithm and confined OMP algorithm, there exists a non-negligible gap between theory and practice. The reasons are as follows.
  • We first consider the lower bound on the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$. The scalings of $\lVert\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}\boldsymbol{x}^{\prime}_{\Lambda^{(k)}}\rVert_{\infty}$ and $\lVert\boldsymbol{A}_{\Omega^{c}}^{T}\boldsymbol{A}_{\Omega}\boldsymbol{x}^{\prime}_{\Lambda^{(k)}}\rVert_{\infty}$ may lead to the gap between the two sides of Eq. 20. Furthermore, the bound on $\lambda_{min}$ derived in Corollary D.5 is loose for a large sparsity $K$, which causes that the bound on $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$ also being loose.
  • We further consider the lower bound on the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$. This bound given in Eq. 10 is actually the exact recovery probability conditioned on the event $\{|\Gamma|=K\}$. Thus, the gap between the analytical curve and experimental curve is equal to the exact recovery probability conditioned on the event $\{|\Gamma|>K\}$. This conditional probability derived in Appendix F is loose for the same reasons as in the case of bound for the OMP algorithm.
Remark: In [35] and [40], lower bounds on the exact recovery probability of OMP algorithm over Gaussian matrices also are loose. However, these theoretical results may provide a guide to determine whether the greedy algorithms are appropriate for reconstruction of sparse signals [35]. Otherwise, another sparse recovery algorithm is considered instead.
Figure 7: Empirical and theoretical recovery performances of confined OMP as a function of column degree $d$.
4.4 Optimization of column degree $d$
As shown in Eq. 10, the lower bound on $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ is related to the column degree $d$ of $\boldsymbol{A}$. We expect to optimize the recovery performance of confined OMP algorithm by optimizing the lower bound on $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$. In other words, the optimized target is to increase the probability $\mathbb{P}\{|\Gamma|=K\}$. In the following experimental simulations, the column degree $d$ of a measurement matrix $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ satisfies $d>\log m$ such that the probability that the least squares have a unique solution is approximately equal to 1. The experimental and theoretical simulations are based on flat signals with $K=10$ and $K=5$, respectively. It can be seen from Fig. 7 that the trend of theoretical curve is almost the same with that of experimental curve. Both of them achieve the best recovery performance when $d=12$.
5 Conclusion and future research problem
This paper proposed a variant of OMP algorithm, referred to as confined OMP algorithm, to recover a class of sparse signals. We proved that the support of $\boldsymbol{x}$ is contained in the confined set $\Gamma$ if the signal $\boldsymbol{x}$ is defined in Definition 2.7. We further presented the expectation of $|\Gamma|$ to show that much redundancy of $\boldsymbol{A}$ can be removed, resulting in a improvement of the proposed algorithm in terms of identification efficiency. We also developed lower bounds on the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$ and $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ over sparse random combinatorial matrices. Finally, experimental results showed that confined OMP algorithm achieves a recovery performance gain and a reduction in complexity compared with OMP algorithm.
There are several future research problems arising from the confined OMP algorithm:
  • As shown in [13], $K<spark(\boldsymbol{A})/2$ is a sufficient and necessary condition for the perfect recovery of a $K$-sparse signal with Eq. 2, where $spark(\boldsymbol{A})$ is the spark of $\boldsymbol{A}$ and defined as $spark(\boldsymbol{A}):=\min\{\lVert\boldsymbol{x}\rVert_{0}:\boldsymbol{A}\boldsymbol{x}=\boldsymbol{0},\boldsymbol{x}\neq\boldsymbol{0}\}$. It is advisable to directly perform $\hat{\boldsymbol{x}}_{\Gamma}=\underset{\boldsymbol{x}\in\mathbb{R}^{|\Gamma|}}{\operatorname*{arg\,min}}\lVert\boldsymbol{y}-\boldsymbol{A}_{\Gamma}\boldsymbol{x}\rVert_{2}$ if $|\Gamma|<spark(\boldsymbol{A})$. In this case, only one least squares needs to be performed. However, the main challenge is to determine the spark of $\boldsymbol{A}$. Fortunately, Theorem D.1 implies that, for a sufficiently large $m$, the spark of a sparse random combinatorial matrix is equal to $m+1$ with probability $1-o(1)$. Our future work will focus on addressing this problem.
  • In many applications, the sparsity $K$ is unknown in practice. Fortunately, as shown in Eq. 5, the expectation of $\nu^{(K)}$ is related to the sparsity $K$ if the signal is defined in Definition 2.7. The sparsity $K$ is predictable according to the sparsity of $\boldsymbol{y}$. It may be interesting to design a blind recovery algorithm that can achieve the same recovery performance as a non-blind recovery algorithm with a moderate increase in complexity.
  • As shown in Corollary D.5, the minimum eigenvalue $\lambda_{min}$ of $\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}$ ranges from $1-r_{m}$ to $1+r_{m}$, where $r_{m}\in\{0,1/d,\cdots,(K-1)d\}$. Since $\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}$ is considered as a positive definite matrix, we have $\lambda_{min}>0$. As $K$ increases, the lower bound on $\lambda_{min}$ becomes less tight, resulting in a loose lower bound on $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$. It may be challenging to give a tight bound on $\lambda_{min}$ or $\sigma_{min}$ in the future.
Appendix A The proof of Theorem 2.2
Assume that the sparse signal $\boldsymbol{x}$ has exactly $K$ non-zero elements. The following Lemma is useful for the proof of Theorem 2.2.
Lemma A.1.
Suppose that the $K$ non-zero components of $\boldsymbol{x}$ in Eq. 1, denoted without loss of generality by $X_{1},X_{2},\cdots,X_{K}$, are i.i.d., with the same CDF $F_{X}(x)$ and PDF $f_{X}(x)$. Given $|y_{i}|\leq\epsilon$, for $\ell=1,2,\cdots,K$ and $i\in\mathcal{E}$, the probability $\mathbb{P}\{|\Gamma_{i}^{c}\cap\Omega|=\ell\}$ is given by
(11) $\mathbb{P}\{|\Gamma_{i}^{c}\cap\Omega|=\ell\}=\binom{K}{\ell}\left(F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\right).$
Proof A.2.
The event $\left\{|\Gamma_{i}^{c}\cap\Omega|=\ell\right\}$ is equivalent to the event that the sum of $\ell$ (out of $K$) non-zero components is less than or equal to $\epsilon$. Without loss of generality, we assume that $Y=X_{1}+X_{2}+\cdots+X_{\ell}$. Consider the simple case that $\ell=2$, the CDF of $Y=X_{1}+X_{2}$ is calculated as
[…]
$F_{Y}(y)=(\underbrace{F_{X}*F_{X}*\cdots*F_{X}}_{\ell\text{ times}})(y)=F_{X}^{*\ell}(y).$
Once the CDF $F_{Y}(y)$ is obtained, one can calculate the probability of the event $\{|\Gamma_{i}^{c}\cap\Omega|=\ell\}$:
$\mathbb{P}\{|\Gamma_{i}^{c}\cap\Omega|=\ell\}=\binom{K}{\ell}\left(F_{Y}(\epsilon)-F_{Y}(-\epsilon)\right)=\binom{K}{\ell}(F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)).$
In the following, we present the proof of Theorem 2.2.
Proof A.3.
The support of $\boldsymbol{x}$ is a subset of $\Gamma$ if and only if the support of $\boldsymbol{x}$ is disjoint from the set $\Gamma^{c}$. In other words, the event $\{\Omega\subseteq\Gamma\}$ is equivalent to the event $\{|\Gamma^{c}\cap\Omega|=0\}$. Further, the event $\{|\Gamma^{c}\cap\Omega|=0\}$ is equivalent to the event that $\{|\Gamma_{i}^{c}\cap\Omega|=0\}$ for all $i\in\mathcal{E}$. Thus, we have
$\mathbb{P}\{\Omega\subseteq\Gamma\}$ =$=\mathbb{P}\left\{|\Gamma^{c}\cap\Omega|=0\right\}$
=$=\mathbb{P}\left\{\Bigg{|}\bigcup_{i\in\mathcal{E}}\Gamma_{i}^{c}\cap\Omega\Bigg{|}=0\right\}$
$=\prod_{i\in\mathcal{E}}\mathbb{P}\{|\Gamma_{i}^{c}\cap\Omega|=0\}$
$=\prod_{i\in\mathcal{E}}\big{(}1-\mathbb{P}\{|\Gamma_{i}^{c}\cap\Omega|\neq 0\}\big{)}$
(12a) $=\left(1-\sum_{\ell=1}^{K}\binom{K}{\ell}\Big{(}F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\Big{)}\right)^{|\mathcal{E}|},$
where Eq. 12a follows from Lemma A.1.
Appendix B The proof of Lemma 3.1
Proof B.1.
It is known that $\boldsymbol{y}$ is a linear combination of $K$ columns of $\boldsymbol{A}$. For $K=1$, the value of $\nu^{(1)}$ is always equal to $d$ since the degree of each column of $\boldsymbol{A}$ is $d$. As a result, we have
$\mathbb{P}\{\nu^{(1)}=d\}=1.$
For $k=2,3,\cdots,K$, the value of $\nu^{(k)}$ ranges from $d$ to $kd$. Here, the value of $\nu^{(k)}$ is equivalent to the number of non-zero elements for a vector resulting from the element-wise OR operation applied to $k$ columns of $\boldsymbol{A}$. It is hard to calculate the probability of $\nu^{(k)}$ directly since it is related to $k$ columns of $\boldsymbol{A}$. Fortunately, it can be modeled as a Markov model. Specifically, by the law of total probability, the probability $\mathbb{P}(\nu^{(k)})$ is
(13) $\mathbb{P}(\nu^{(k)})=\sum_{\nu^{(k-1)}}\mathbb{P}(\nu^{(k)}|\nu^{(k-1)})\mathbb{P}(\nu^{(k-1)}).$
The transition probability $\mathbb{P}(\nu^{(k)}|\nu^{(k-1)})$ only depends on $\nu^{(k-1)}$ and $\nu^{(k)}$. This is equivalent to a linear combination of a vector with $\nu^{(k-1)}$ non-zero elements and any column of $\boldsymbol{A}$. If $\nu^{(k-1)}<\max\{\nu^{(k)}-d,d\}$ and $\nu^{(k-1)}>\min\{\nu^{(k)},(k-1)d\}$, the transition probability $\mathbb{P}(\nu^{(k)}|\nu^{(k-1)})$ is obviously 0. Otherwise, the transition probability is
(14) $\mathbb{P}(\nu^{(k)}|\nu^{(k-1)})=\frac{\binom{\nu^{(k-1)}}{\nu^{(k)}-\nu^{(k-1)}}\binom{m-\nu^{(k-1)}}{d-\nu^{(k)}+\nu^{(k-1)}}}{\binom{m}{d}},$
where $\nu^{(k)}\in\{d,d+1,\cdots,kd\}$ and $\nu^{(k-1)}\in\{\max\{\nu^{(k)}-d,d\},\max\{\nu^{(k)}-d,d\}+1,\cdots,\min\{\nu^{(k)},(k-1)d\}\}$. Here,  Eq. 14 represents the probability that any column of $\boldsymbol{A}$ shares exactly $\nu^{(k)}-\nu^{(k-1)}$ non-zero positions with a vector having $\nu^{(k-1)}$ non-zero elements.
The probability $\mathbb{P}(\nu^{(k-1)})$ can be calculated by Eq. 13 recursively. As a consequence, we get Eq. 4.
Appendix C The proof of Theorem 3.4
Proof C.1.
As shown in Theorem 2.2, the support of $\boldsymbol{x}$ is a subset of the confined set $\Gamma$ with probability 1. Thus, there are at least $K$ out of $n$ columns whose indices are in $\Gamma$. That is, we have $\mathbf{E}[|\Gamma|]\geq K$.
(15) $\mathbb{P}\left\{j\in\Gamma\backslash\Omega\Big{|}\nu^{(K)}=\upsilon\right\}=\frac{\binom{\upsilon}{d}}{\binom{m}{d}}.$
Here, (15) indicates the probability that the column $\boldsymbol{A}_{j}$ shares exactly $d$ non-zero positions with $\boldsymbol{y}$ having $\upsilon$ non-zero elements. If this is not the case, then $j\notin\Gamma$ since the $i$-th element of $\boldsymbol{A}_{j}$ must be 0 for $i\in\mathcal{E}$.
Following from Lemma 3.1, the value of $\upsilon$ ranges from $d$ to $Kd$. Thus, by the law of total probability, we have
(16) $\begin{split}\mathbb{P}\left\{j\in\Gamma\backslash\Omega\right\}&=\sum_{\upsilon=d}^{Kd}\mathbb{P}\left\{j\in\Gamma\backslash\Omega\Big{|}\nu^{(K)}=\upsilon\right\}\mathbb{P}\left\{\nu^{(K)}=\upsilon\right\}\\ &=\sum_{\upsilon=d}^{Kd}\frac{\binom{\upsilon}{d}}{\binom{m}{d}}\mathbb{P}\left\{\nu^{(K)}=\upsilon\right\},\end{split}$
where $\mathbb{P}\left\{\nu^{(K)}=\upsilon\right\}$ is given in Lemma 3.1.
In summary, there are $K$ columns whose indices are in $\Omega\subseteq\Gamma$. For the remaining $n-K$ column indices, each of them belongs to $\Gamma\backslash\Omega$ with probability given in Eq. 16. As a result, we get $\mathbf{E}[|\Gamma|]$ shown in Eq. 6.
Appendix D The proof of Theorem 3.6
There are two Corollaries used in the proof of Theorem 3.6. These two Corollaries are obtained from two different Theorems, which will be proved below respectively.
Theorem D.1 (Theorem 1.2 in [16]).
Fix $\varepsilon>0$, and let $d=d(m)$ be any function of $m$ satisfying $\min\{d,m-d\}\geq(1+\varepsilon)\log m$. Then, for an $m\times m$ random combinatorial matrix $\boldsymbol{Q}$ with independent rows, where each row is chosen uniformly among the vectors with $d$ ones, we have
Corollary D.2.
Fix $\varepsilon>0$, and let $d=d(m)$ be any function of $m$ satisfying $(1+\varepsilon)\log m\leq d\leq m/2$. Then, for an $m\times K$ random combinatorial matrix $\boldsymbol{A}_{\Omega}$ with $d$ ones per column and a sufficiently large $m$, we have
(17) $\mathbb{P}\left\{\det\left(\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}\right)\neq 0\right\}=1-o(1),$
where $\det(\cdot)$ represents the determinant of argument.
Proof D.3.
According to Theorem D.1, we know that the $m\times m$ random combinatorial matrix $\boldsymbol{Q}$ is nonsingular with probability $1-o(1)$ for a sufficiently large $m$. In other words, $m$ columns of $\boldsymbol{Q}^{T}$ are linearly independent with probability $1-o(1)$ for a sufficiently large $m$. For the $m\times K$ random combinatorial matrix $\boldsymbol{A}_{\Omega}$ with a sufficiently large $m$, $K$ columns of $\boldsymbol{A}_{\Omega}$ are obviously linearly independent with probability $1-o(1)$. Furthermore, it is known that the rank of $\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}$ is equal to that of $\boldsymbol{A}_{\Omega}$. Thus, we get Eq. 17.
Remark: In general, the sparsity $K$ is much smaller than $m$. Thus, as the increase of $m$, the probability $\mathbb{P}\{\det\left(\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}\right)\neq 0\}$ converges to 1 more rapidly than the probability $\mathbb{P}\{\det(\boldsymbol{Q})\neq 0\}$ shown in Theorem D.1.
Let $\boldsymbol{M}$ be an $m\times m$ matrix and $r_{i}=\sum_{j=1,j\neq i}^{m}|M_{i,j}|$ be the sum of the absolute values of the non-diagonal elements in the $i$-th row of $\boldsymbol{M}$ for $i=1,2,\cdots,m$. Define a circle centered at $M_{i,i}$ with a radius of $r_{i}$ as a Gershgorin disc $D(M_{i,i},r_{i})$.
Theorem D.4 (Gershgorin’s circle Theorem [21] ).
Every eigenvalues of $\boldsymbol{M}$ lies in one of Gershgorin discs $D(M_{i,i},r_{i})$.
Corollary D.5.
Let $\boldsymbol{G}=\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}$. The minimum eigenvalue $\lambda_{min}$ of $\boldsymbol{G}$ lies in the Gershgorin disc $D(1,r_{m})$, where $r_{m}=\underset{i\in\Omega}{\max}\sum_{j\in\Omega,j\neq i}\mu_{i,j}$.
Proof D.6.
Without loss of generality, we assume that $\Omega=\{1,2,\cdots,|\Omega|\}$. The symmetric matrix $\boldsymbol{G}$ is given by
$\boldsymbol{G}=\boldsymbol{A}_{\Omega}^{T}\boldsymbol{A}_{\Omega}=\begin{pmatrix}1&\mu_{1,2}&\mu_{1,3}&\cdots&\mu_{1,|\Omega|}\\ \mu_{2,1}&1&\mu_{2,3}&\cdots&\mu_{2,|\Omega|}\\ \mu_{3,1}&\mu_{3,2}&1&\cdots&\mu_{3,|\Omega|}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mu_{|\Omega|,1}&\mu_{|\Omega|,2}&\mu_{|\Omega|,3}&\cdots&1\end{pmatrix}.$
It can be seen that the centers of all Gershgorin discs are the same. Thus, all eigenvalues of $\boldsymbol{G}$ lie in the largest Gershgorin disc, i.e., $D(1,r_{m})$.
Assume that each column of $\boldsymbol{A}$ is normalized. The proof of Theorem 3.6 is presented as follows.
Proof D.7.
Define $\mathbb{D}$ as the event $\left\{\det\left(\boldsymbol{A}_{\Lambda^{(K)}}^{T}\boldsymbol{A}_{\Lambda^{(K)}}\right)\neq 0\right\}$. To ensure the $K$-sparse signal $\boldsymbol{x}$ can be exactly recovered by the OMP algorithm in $K$ iterations, the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}$ is equivalent to
(18) $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}=\prod_{k=1}^{K}\mathbb{P}\left\{\lVert\boldsymbol{A}_{\Omega}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}>\lVert\boldsymbol{A}_{\Omega^{c}}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}\Big{|}\mathbb{D}\right\}\mathbb{P}\left\{\mathbb{D}\right\}.$
During the identification step of OMP algorithm, the algorithm picks a “true” column from $\boldsymbol{A}_{\Omega}$ whenever $\lVert\boldsymbol{A}_{\Omega}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}>\lVert\boldsymbol{A}_{\Omega^{c}}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}$. These chosen columns determine which elements of $\boldsymbol{x}$ are non-zero. The values of these elements of $\boldsymbol{x}$ are determined by solving a least-squares problem, which has a unique solution if $\boldsymbol{A}_{\Lambda^{(K)}}^{T}\boldsymbol{A}_{\Lambda^{(K)}}$ has full rank, i.e., $\det\left(\boldsymbol{A}_{\Lambda^{(K)}}^{T}\boldsymbol{A}_{\Lambda^{(K)}}\right)\neq 0$. Based on Corollary D.2, the probability $\mathbb{P}\left\{\mathbb{D}\right\}=1-o(1)$ if the column degree $d$ satisfies $(1+\varepsilon)\log m\leq d\leq m/2$ for a constant $\varepsilon>0$ and a sufficiently large $m$. Thus, we have
(19) $\mathbb{P}\{\mathbb{S}_{\mathrm{omp}}\}\geq\prod_{k=1}^{K}\mathbb{P}\left\{\lVert\boldsymbol{A}_{\Omega}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}>\lVert\boldsymbol{A}_{\Omega^{c}}^{T}\boldsymbol{r}^{(k-1)}\rVert_{\infty}\Big{|}\mathbb{D}\right\}-o(1).$
In Eq. 19, the first term is based on the assumption that the OMP algorithm has picked $k$ “true” columns from $\boldsymbol{A}_{\Omega}$ at the $(k+1)$-th iteration. Thus, the residual $\boldsymbol{r}^{(k-1)}$ can be rewritten as [38]
$\boldsymbol{r}^{(k-1)}=\boldsymbol{y}-\boldsymbol{A}_{\Lambda^{(k)}}^{T}\hat{\boldsymbol{x}}_{\Lambda^{(k)}},$
where the estimated support set $\Lambda^{(k)}\subseteq\Omega$. Since $\boldsymbol{y}=\boldsymbol{A}_{\Omega}\boldsymbol{x}_{\Omega}$ and $\boldsymbol{A}_{\Lambda^{(k)}}$ is a submatrix of $\boldsymbol{A}_{\Omega}$, the residual $\boldsymbol{r}^{(k-1)}$ also can be expressed as a linear combination of columns of $\boldsymbol{A}_{\Omega}$ [38], resulting in
[…]

### v2

@@ Appendix C The Proof of Theorem 2
where $\mathbb{P}\left\{\nu^{(K)}=\upsilon\right\}$ is given in (4) and $\mathbf{E}_{\upsilon}[\cdot]$ is given in (5).
In summary, there are $K$ columns whose indices are in $\Omega\subseteq\Gamma$. For the remaining nKn-K column indices, each of them belongs to $\Gamma\backslash\Omega$ with probability $\mathbf{E}_{\upsilon}\left[\frac{\binom{\nu^{(K)}}{d}}{\binom{m}{d}}\right$]. As a result, we get $\mathbf{E}[|\Gamma|]$ shown in (7). ∎
Appendix D The proof of Theorem 4
Proof:
By the law of total probability, we have
$\mathbb{P}$ $\{\mathbb{S}_{\mathrm{comp}}\}=\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|=K\right\}\mathbb{P}\left\{|\Gamma|=K\right\}$
$+\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|>K\right\}\mathbb{P}\left\{|\Gamma|>K\right\}$
$=1\cdot\mathbb{P}\left\{|\Gamma|=K\right\}+\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|>K\right\}\mathbb{P}\left\{|\Gamma|>K\right\}$ (34a)
$\geq\mathbb{P}\left\{|\Gamma|=K\right\}.$ (34b)
As shown in Algorithm 2, the identification is already done if $|\Gamma|=K$. Thus, the probability $\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|=K\right\}$ is reduced to the probability that the least squares has a unique solution. Since $\boldsymbol{A}$ satisfies $spark(\boldsymbol{A})>K$, any $K$ columns of $\boldsymbol{A}$ are linearly independent. In other words, $\boldsymbol{A}_{\Lambda^{(K)}}^{T}\boldsymbol{A}_{\Lambda^{(K)}}$ is of full rank. Thus, we have $\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|=K\right\}=1$. In (34a), calculating $\mathbb{P}\left\{\mathbb{S}_{\mathrm{comp}}\big||\Gamma|>K\right\}$ is complicated. Hence, we only consider the contribution of the first term of (34a) for simplicity, resulting in (34b).
Given any $j\in\Omega^{c}$, the probability $\mathbb{P}\left\{|\Gamma|=K\right\}$ is equivalent to the probability $\mathbb{P}\{j\notin\Gamma\}^{|\Omega^{c}|}$. It is known that the value of $\upsilon$ ranges from $d$ to $\min\{Kd,m\}$. Given $\nu^{(K)}=\upsilon$, the probability $\mathbb{P}\left\{j\in\Gamma\backslash\Omega\big|\nu^{(K)}=\upsilon\right\}$ is given in (32). Then, we have
$\mathbb{P}\{|\Gamma|=K\}$ $=\mathbb{P}\{j\notin\Gamma\}^{|\Omega^{c}|}$
$=\left(1-\mathbb{P}\left\{j\in\Gamma\backslash\Omega\right\}\right)^{n-K}$ (35a)
$=\sum_{\upsilon=d}^{\min\{Kd,m\}}\left(1-\frac{\binom{\upsilon}{d}}{\binom{m}{d}}\right)^{n-K}\mathbb{P}\{\nu^{(K)}=\upsilon\},$ (35b)
$=\mathbf{E}_{\upsilon}\left[\left(1-\frac{\binom{\nu^{(K)}}{d}}{\binom{m}{d}}\right)^{n-K}\right],$ (35c)
where $\mathbb{P}\{\nu^{(K)}=\upsilon\}$ is given in (4) and $\mathbf{E}_{\upsilon}[\cdot]$ is given in (5). Thus, we get (9). ∎
Appendix E The proof of Corollary 4
Proof:
Since $K\leq m/d$, the value of $\nu^{(K)}$ ranges from $d$ to $Kd$. With the Jensen’s inequality, the probability $\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ is lower bounded by
$\mathbb{P}\{\mathbb{S}_{\mathrm{comp}}\}$ $\geq\left(1-\mathbf{E}_{\upsilon}\left[\frac{\binom{\nu^{(K)}}{d}}{\binom{m}{d}}\right]\right)^{n-K}$ (36a)
$\geq\left(1-\frac{\binom{Kd}{d}}{\binom{m}{d}}\right)^{n-K}$ (36b)
$=\left(1-\prod_{z=0}^{d-1}\frac{Kd-z}{m-z}\right)^{n-K}$ (36c)
$\geq\left(1-\left(\frac{Kd}{m}\right)^{d}\right)^{n-K}=\bar{\pi}_{K}.$ (36d)
With all parameter in $\bar{\pi}_{K}$ fixed except for $d$, we derive the maximum value of $\bar{\pi}_{K}(d)$ with respect to $d$. Let $h(d)=\left(\frac{Kd}{m}\right)^{d}$. We first need to perform differentiation on $h(d)$. Before that, we respectively take the logarithm of both sides, i.e.,
$\ln h(d)=d\ln\left(\frac{Kd}{m}\right).$ (37)
Then, we have
$\left(\ln h(d)\right)^{\prime}$ $=\left(d\ln\left(\frac{Kd}{m}\right)\right)^{\prime},$ (38a)
$\frac{h^{\prime}(d)}{h(d)}$ $=\ln\left(\frac{Kd}{m}\right)+d\cdot\frac{m}{Kd}\cdot\frac{K}{m},$ (38b)
$h^{\prime}(d)$ $=h(d)\cdot\left(\ln\left(\frac{Kd}{m}\right)+1\right).$ (38c)
Now, we perform differentiation on $\bar{\pi}_{K}(d)$. Similarly, we have
$\ln\bar{\pi}_{K}(d)$ $=(n-K)\ln\left(1-h(d)\right),$ (39a)
$\left(\ln\bar{\pi}_{K}(d)\right)^{\prime}$ $=\left((n-K)\ln\left(1-h(d)\right)\right)^{\prime},$ (39b)
$\bar{\pi}_{K}^{\prime}(d)$ $=-\bar{\pi}_{K}(d)\cdot(n-K)\frac{h^{\prime}(d)}{1-h(d)},$ (39c)
$\bar{\pi}_{K}^{\prime}(d)$ $=-\bar{\pi}_{K}(d)\cdot(n-K)\frac{h(d)}{1-h(d)}\left(\ln\left(\frac{Kd}{m}\right)+1\right).$ (39d)
Let $\bar{\pi}_{K}^{\prime}(d)=0$. Since $\bar{\pi}_{K}(d)>0$, $h(d)>0$, and nK>0n-K>0, $\bar{\pi}_{K}^{\prime}(d)=0$ is equivalent to $\ln\left(\frac{Kd}{m}\right)+1=0$. A simple derivation can yield that $\bar{\pi}_{K}(d)$ reaches its maximum value when $d=\frac{m}{K\cdot\mathrm{e}}$. By substituting $d=\frac{m}{K\cdot\mathrm{e}}$ into $\bar{\pi}_{K}(d)$, we get $\bar{\pi}_{K}^{*}=\left(1-\mathrm{e}^{-\frac{m}{K\cdot\mathrm{e}}}\right)^{n-K}$. ∎
Appendix F The proof of Theorem 5
Proof:
[…]